OpenStudy (anonymous):

224 x 37over56 x 9over16 =

OpenStudy (espex):

Is your question \[224 \times \frac{37}{56} \times \frac{9}{16}=?\]

OpenStudy (anonymous):

yes

OpenStudy (espex):

Since there are no variables here, you can multiply the numerators together, the denominators together, and finally divide the two products.

OpenStudy (anonymous):

how about the 224

OpenStudy (espex):

Any whole number can be rewritten as a fraction like this: \[\frac{224}{1}\]